# Breaking the Bulkhead: Demystifying Cross-Namespace Reference Vulnerabilities in Kubernetes Operators

### v1

@@ 5.2.2. RQ2: What Resources Can Be Cross-Namespace Referenced By Attackers?
To understand the attack surface exposed by insecure cross-namespace references, we investigate the types of resources that Operators allow attackers to reference across namespace boundaries. For each resource type, we count the number of Operators that insecurely reference it and analyze which types are most frequently involved in such behavior.
@@
Among namespace-scoped resources, the most commonly insecurely referenced types (listed in Table 3) are Secret (referenced by 102 Operators), ConfigMap (29 Operators), and Deployment (29 Operators). In Kubernetes, Secrets store highly sensitive data such as API keys, credentials, and TLS certificates. ConfigMaps often contain important application configurations that control applications behavior, like API endpoints and performance arguments. Deployments define and manage the application workloads by controlling replica sets and pods.
@@
For cluster-scoped resources, the most common insecurely referenced types are Namespace (62 Operators), ClusterRoleBinding (40 Operators), and ClusterRole (26 Operators). In Kubernetes, Namespaces are resources that define the namespace in a Kubernetes cluster. ClusterRoles and ClusterRoleBindings define and grant cluster-level permissions that apply in the whole cluster.
Figure 7. Reference of Built-In and Custom Resources
@@
We also distinguish between insecure references to built-in Kubernetes resources and custom resources. To this end, we aggregated the type-#Operator result above based on built-in resource type or custom resource type. For insecure namespace-scoped references, 279 cases involved built-in resources and 182 involved custom resources. For insecure cluster-scoped references, 211 targeted built-in resources and 87 involved custom resources. These results indicate that insecure references can affect both built-in resources and custom resources. And the insecure references are more commonly associated with Kubernetes built-in resources.
5.2.3. RQ3: What Can Attacker Do Towards Cross-Namespace Referenced Resources?
@@
Figure 8. Verbs Used By # Operators Towards Insecurely Referenced Resources
@@
To understand the potential impact of insecure cross-namespace references, we analyze the operations (Kubernetes API Verbs) that Operators perform on the referenced resources. For each identified insecurely referenced resource, we extract its related verbs. We then count how many Operators apply each verb to each insecurely referenced resource type. The result is illustrated in Figure 8.
@@
For insecurely referenced namespace-scoped resources, the top three most common verbs are Get (used by 188 Operators), Create (74 Operators), and Update (61 Operators), with Get being the most prevalent. In Kubernetes, Get retrieves the current state of a resource, Create instantiates a new resource, and Update modifies an existing resource. The predominance of the Get operation indicates that a large number of vulnerable Operators retrieve data from resources in other namespaces based on attacker-controlled inputs, exposing unauthorized data to attackers.
@@
Figure 9. Verbs of Major Resources
Table 4. Major Insecurely Used Verb-Resource Pairs
@@
To further understand the practical implications, we investigate the verb-resource pairs (i.e., combinations of verbs and insecurely referenced resource types) to determine which verbs Operators typically perform on specific insecurely referenced resource types. The result is illustrated in Table 4 and Figure 9.
@@
For insecurely referenced namespace-scoped resource types, the most prevalent pairs are Get-Secret (used by 97 Operators), Get-ConfigMap (25 Operators), and Get-Deployment (25 Operators). These indicate that a substantial number of Operators can be exploited to read data from Secrets, configuration files, or application deployments in attackers’ unauthorized namespaces, causing information exposure.
@@
For insecurely referenced cluster-scoped resource types, the most prevalent pairs are Get-Namespace (used by 51 Operators), Create-ClusterRoleBinding (33 Operators), and Create-Namespace (30 Operators). These patterns suggest that Operators may be exploited to reveal other namespaces in the cluster, assign cluster-wide permissions, or provision new namespaces, leading to privilege elevation.
@@
Together, these findings highlight that insecure cross-namespace references are not only present but often tied to high-impact operations on sensitive or privileged Kubernetes resources.
5.3. RQ4: Case Study
We responsibly disclosed vulnerabilities to affected vendors. The case study is hidden due to embargoes.
6. Mitigation
@@
During our inspection of vulnerable Operators, we realized that Operators aim to simplify user operations as much as possible, thus may embed cross-namespace reference functionality to spare users the repetitive, manual task of duplicating resources like Secrets across namespaces—a practice necessitated by Kubernetes’ namespace isolation. However, this convenience can inadvertently introduce vulnerabilities if not properly implemented, as attackers may exploit such ”helpful” behavior to perform unauthorized cross-namespace actions and privilege escalations. Thus, we suggest the following mitigations to eliminate the insecure cross-namespace reference vulnerability.
@@
Scope Alignment. Developers should ensure that the declared scope of resources accurately reflects the scope of their operational effect. If a resource is defined as namespace-scoped but its process logic performs actions across multiple namespaces at the cluster level, it creates a dangerous mismatch between the resource’s access control boundary and its actual impact. In such cases, the resource should be explicitly declared as cluster-scoped, ensuring that only privileged users can create or manipulate it.
Avoid Uncessary Cross-Namespace References. Operator developers should avoid supporting unnecessary cross-namespace references in Custom Resource specifications. Whenever possible, references to other resources, such as Secret and ConfigMap, should be limited to the same namespace as the Custom Resource itself. If cross-namespace references are truly required, it is suggested to receive such requests from cluster-scoped resources, ensuring users have enough permissions and no privilege escalation occurs.
Minimize User Permissions. While an Operator may legitimately require elevated permissions, such as reading Secrets across namespaces, to fulfill its responsibilities, these privileges often cannot be further minimized without breaking functionality. However, permission minimization can still be enforced at the user side. In multi-tenant environments, cluster administrators should minimize user permissions to prevent low-privileged users from exploiting Operators. Role-based access control (RBAC) should be configured so that only trusted users or service accounts can interact with sensitive resources. By carefully controlling who can invoke the Operator, administrators can prevent tenants from abusing it to perform unauthorized cross-namespace or cluster-level operations.
Future Work: Finer Access Control. A fundamental limitation in the current Kubernetes architecture is that Operators typically lack visibility into who initiated a resource request. As future work, we aim to design and implement a mechanism for Operators to authenticate and verify the identity of the request user associated with a resource event. This capability would allow Operators to enforce that privileged operations are only performed on behalf of users with appropriate permissions, strengthening the security of Kubernetes Operators in multi-tenant environments.
7. Related Work
@@
Kubernetes Operators and Controllers. Existing research on Kubernetes operators and controllers focuses on functional bugs instead of security vulnerabilities. Gu et al. (2023) proposes an automatic end-to-end testing technique for validating the operational correctness of Kubernetes Operators. Acto continuously generates desired state declarations and verifies whether the Operator correctly reconciles the system to those states. Sun et al. (2022) presents an automatic reliability testing framework for cluster-management controllers. By injecting faults, Sieve uncovers deep semantic bugs by observing how controllers behave under fault conditions they are expected to tolerate. Sun et al. (2024) presents the formal verification framework for Kubernetes controllers via TLA-style temporal reasoning, validating whether controllers eventually bring the cluster to the desired state and maintain it. Liu et al. (2024) verifies Kubernetes controllers and their configurations by modeling controller behaviors and checking for violations of user-defined intent properties using model checking. It detects issues like imbalance and lifecycle bugs, focusing on functional correctness. Xu et al. (2024) systematically summarizes historical functional bugs of Operator.
@@
To the best of our knowledge, our work is the first comprehensive study on Kubernetes Operator security.
@@
Kubernetes Security. In terms of attack and exploitation techniques, MITRE (mit, 2025) and Microsoft (msm, 2025) summarize tactics to compromise containers and container orchestration systems like Kubernetes. Pecka et al. (2022) investigates privilege escalation scenarios for DevOps pipelines on Kubernetes. He et al. (2023) presents cross-container attacks on Kubernetes with eBPF. Spahn et al. (2023) sets up honeypots on Kubernetes and analyzes the attacks towards containers and container orchestration systems. Shringarputale et al. (2020) presents a co-residency attack towards container orchestration systems. Zeng et al. (2023) comprehensively analyzes 30 vulnerabilities in Kubernetes stacks. However, these attacks have not included the security issues brought about by Kubernetes operators.
@@
Kubernetes offers extensive configuration options for managing applications, including access controls and specifying security contexts. Any misconfigurations can lead to severe security vulnerabilities. Thus, another theme of Kubernetes security research is eliminating misconfiguration. Islam Shamim et al. (2020) systematically identifies best practices to secure Kubernetes clusters. Shamim et al. (2025) conducts an empirical study and reveals the disconnection between Kubernetes configuration recommendations and real-world practices. Rahman et al. (2023) designs static analysis tools and conducts a large-scale empirical study on Kubernetes manifests, revealing the landscape of misconfiguration. Ul Haque et al. (2022) leverages knowledge graphs to detect and mitigate Kubernetes misconfiguration. Shamim (2021) explores the risk of misconfiguration when violating Kubernetes security best practices. Recent work Yang et al. (2023) identifies the security risk of excessive Kubernetes RBAC permissions, which may lead to whole cluster takeover. Gu et al. (2025) follows up the research and designs systems to automatically minimize RBAC permissions. The industry also presents numerous tools for Kubernetes security, including Trivy (tri, 2025), Kubescape (kub, 2025d), KubeSec (kub, 2025a), KubeArmor (kub, 2025b), Open Policy Agent (opa, 2025), and Kyverno (kyv, 2025), providing functions like misconfiguration detection and runtime policy enforcement.
@@
While the existing works try to address misconfiguration and thus achieve the Principle of Least Privilege (PoLP) for applications, the vulnerability we present is not simply misconfigurations or violations of PoLP. They arise from inherent flaws in how Operators process user-controlled resources. These vulnerabilities exist even when the permissions of Operators are minimal in the current Kubernetes architecture, highlighting a deeper design-level security gap in the Operator model itself.
8. Conclusion

### v2

@@ V-B2 RQ2: What Resources Can Be Cross-Namespace Referenced By Attackers?
To understand the attack surface exposed by insecure cross-ns references, we investigate the types of resources that Operators allow attackers to reference across namespaces. For each resource type, we count the number of Operators that insecurely reference it and analyze which types are most frequently involved.
@@
Among ns-scoped resources, the most commonly insecurely referenced types (listed in Table IV) are Secret (referenced by 102 Operators), ConfigMap (29 Operators), and Deployment (29 Operators). In Kubernetes, Secrets store highly sensitive data such as API keys, credentials, and TLS certificates. ConfigMaps often contain important application configurations that control applications’ behavior, like API endpoints and performance arguments. Deployments define and manage the application workloads by controlling replica sets and pods.
@@
For cluster-scoped resources, the most common insecurely referenced types are Namespace (62 Operators), ClusterRoleBinding (40 Operators), and ClusterRole (26 Operators). In Kubernetes, Namespace is the resource that defines a namespace in a Kubernetes cluster. ClusterRoles and ClusterRoleBindings define and grant cluster-level permissions that apply to the whole cluster.
Figure 7: Reference of Built-In and Custom Resources
@@
We also distinguish between insecure references to built-in Kubernetes resources and custom resources. To this end, we aggregated the type-#Operator result above based on built-in resource type or custom resource type. For insecure ns-scoped references, 279 cases involved built-in resources and 182 involved custom resources. For insecure cluster-scoped references, 211 targeted built-in resources and 87 involved custom resources. These results indicate that insecure references can affect both built-in resources and custom resources. And the insecure references are more commonly associated with Kubernetes built-in resources.
V-B3 RQ3: What Can Attackers Do Towards Cross-Namespace Referenced Resources?
@@
Figure 8: Verbs Used By # Operators Towards Insecurely Referenced Resources
@@
To understand the potential impact of insecure cross-ns references, we analyze the operations (Kubernetes API Verbs) that Operators perform on the referenced resources. For each identified insecurely referenced resource, we extract its related verbs. We then count how many Operators apply each verb to each insecurely referenced resource type. The result is illustrated in Figure 8.
@@
For insecurely referenced ns-scoped resources, the top three most common verbs are Get (used by 188 Operators), Create (74 Operators), and Update (61 Operators), with Get being the most prevalent. In Kubernetes, Get retrieves the current state of a resource, Create instantiates a new resource, and Update modifies an existing resource. The predominance of the Get operation indicates that a large number of vulnerable Operators retrieve data from resources in other namespaces based on attacker-controlled inputs, exposing unauthorized data to attackers.
@@
Figure 9: Verbs of Major Insecurely Referenced Resources
TABLE V: Major Insecurely Used Verb-Resource Pairs
@@
To understand the practical implications, we investigate the verb-resource pairs (i.e., combinations of verbs and insecurely referenced resource types) to determine which verbs Operators typically perform on specific insecurely referenced resource types. The result is illustrated in Table V and Figure 9.
@@
For insecurely referenced ns-scoped resource types, the most prevalent pairs are Get-Secret (used by 97 Operators), Get-ConfigMap (25 Operators), and Get-Deployment (25 Operators). These indicate that a substantial number of Operators can be exploited to read data from Secrets, configuration files, or application deployments in attackers’ unauthorized namespaces, causing information exposure.
@@
For insecurely referenced cluster-scoped resource types, the most prevalent pairs are Get-Namespace (used by 51 Operators), Create-ClusterRoleBinding (33 Operators), and Create-Namespace (29 Operators). These patterns suggest that Operators may be exploited to reveal other namespaces in the cluster, assign cluster-wide permissions, or provision new namespaces, leading to privilege elevation.
@@
Together, these findings highlight that insecure cross-ns references are not only present but often tied to high-impact operations on sensitive or privileged Kubernetes resources.
V-C RQ4: Case Study
To validate the practical significance of the identified vulnerabilities, we conducted in-depth inspections of vulnerable Operators. Combined with static analysis and manual exploits, the detailed impacts of their vulnerabilities are identified. We responsibly disclosed our findings to their vendors. By the time of submission, 8 vulnerabilities have been confirmed by the affected vendors as listed in Table VI, where all vulnerable resources are custom resources, as Operators are typically triggered by their custom resources and then reference other resources. 7 CVEs have either been assigned or are currently under processing, reflecting community acknowledgment and real-world relevance of these issues. Notably, even Operators maintained by Red Hat, the inventor of Operators, were confirmed to be vulnerable to insecure cross-ns references, highlighting that the problem is systemic and not limited to less mature projects.
We introduced the case of Grafana/tempo-operator to illustrate how these insecure references impact the real world.
TABLE VI: Vendor-Confirmed Vulnerabilities
Operator Vendor Status Vulnerable Resource Referenced Resource Verb
Operator X 1 ****** Confirmed ****** ****** ******
gateway-operator [konggateway] Kong CVE Assigning AIGateway Secret Get
baremetal-operator Metal3-io CVE-2025-29781 BMCEventSubscription Secret Get
observability-operator [observability] Red Hat CVE-2025-2843 MonitorStack ClusterRoleBinding Create/Update/Delete
gateway-operator [konggateway] Kong CVE Assigning ControlPlane ClusterRoleBinding Create
tempo-operator [tempo] Grafana & Red Hat CVE-2025-2842 TempoStack ClusterRoleBinding Create/Delete
tempo-operator [tempo] Grafana & Red Hat CVE-2025-2786 TempoMonolithic ClusterRoleBinding Create/Delete
ais-k8s NVIDIA CVE-2025-23260 AIStore ClusterRoleBinding Create/Update/Delete
  • 1
    Hidden for Ethical Consideration.
Grafana [cncf], a member of CNCF, is a widely adopted open-source analytics and visualization platform, renowned for transforming complex data into interactive dashboards. Among their products, Grafana Tempo [tempo-intro] stands as a distributed tracing backend for the Grafana Observability Stack, gaining over 4,000 stars on GitHub. The tempo-operator [tempo], developed by both Grafana and Red Hat, is the official solution for deploying and managing Grafana Tempo on Kubernetes clusters. It defines a ns-scoped Custom Resource, TempoStack, for users to deploy Tempo.
Specifically, when enabling JaegerQuery and MonitorTab functions in a TempoStack resource, tempo-operator would create a ClusterRoleBinding, granting cluster-level permissions to a ServiceAccount in the namespace specified by the attacker-controlled TempoStack.Namespace field. An attacker can thus deploy a TempoStack in his namespace, where the JaegerQuery and MonitorTab functions were enabled and TempoMonolithic.Namespace field was set to the name of his authorized namespace. The Operator will then grant cluster-level permissions to a ServiceAccount in his namespace. The attacker can then impersonate the ServiceAccount to gain access to resources in the whole cluster.
In terms of impacts, the operator would grant the cluster-monitoring-view ClusterRole, an OpenShift-specific role enabling access to Prometheus and Thanos APIs, thus allowing the attacker to observe metrics and monitor cluster-wide resource states. We validated the bug and reported it to the vendors. They responded that “We confirmed the vulnerability and will start the process to assign a CVE shortly. ”. CVE-2025-2842 was then assigned to us.
VI Mitigation & Discussion
@@
During our inspection of vulnerable Operators, we realized that Operators aim to simplify user operations as much as possible, thus may embed cross-ns reference functionality to spare users the repetitive, manual task of duplicating resources like Secrets across namespaces—a practice necessitated by Kubernetes’ namespace isolation. However, this convenience can inadvertently introduce vulnerabilities if not properly implemented, as attackers may exploit such helpful behavior to perform unauthorized cross-ns actions and privilege escalations. Thus, we suggest the following mitigations to eliminate the cross-ns reference vulnerability.
Carefully Using Multi-Tenant Kubernetes. Practitioners are discouraged from sharing Kubernetes across untrusted users, ensuring all tenants in multi-tenant Kubernetes do no evil. Developers should eliminate vulnerabilities of applications running on multi-tenant clusters, which may be exploited to gain initial Kubernetes access and affect other tenants.
@@
Scope Alignment. Developers are advised to ensure that the declared scope of resources accurately reflects the scope of their operational effect. If a resource is defined as ns-scoped but its process logic performs actions across multiple namespaces at the cluster level, it creates a dangerous mismatch between the resource’s access control boundary and its actual impact. In such cases, the resource should be explicitly declared as cluster-scoped, ensuring that only privileged users can create or manipulate it.
Mitigate with Kubernetes Admission Control. In cases where cross-ns is necessary or modifying Operators is painful, we designed a mitigation approach based on Validating Admission Webhook [admissioncontrol] to reject harmful requests before proceeding to the original Operators. The mitigation works independently, demands no modification to existing Operators, and can be easily adapted to any project.
Essentially, Validating Admission Webhook is a mechanism to extend the native Kubernetes access control. When a user manipulates a resource, besides RBAC, Kubernetes will also invoke Webhooks to decide whether to allow or reject the request. In addition, Kubernetes provides the SubjectAccessReview [subjectaccessreview] API, which can be used to check whether a user has specific RBAC permissions on a given resource. We leverage these mechanisms to further check whether the user has permissions for referenced resources and inform Kubernetes to reject requests if the user is unauthorized. This fills the gap of the native Kubernetes RBAC, which is not aware of whether a resource has cross-ns references.
Figure 10: Mitigate with Validating Admission Webhook
Figure 10 gives the workflow of our proposed mitigation. The basic flow of a Validating Admission Webhook is that the Kubernetes API Server intercepts a user request, sends an AdmissionReview object containing the raw resource and user identity to the Webhook, and waits for an AdmissionResponse object from the Webhook indicating whether the request should be accepted or rejected. Upon receiving an AdmissionReview from the Kubernetes API Server, the Webhook reads the object and evaluates whether the user has cross-namespace permissions by issuing a SubjectAccessReview query to the API Server. If the result indicates that the user lacks permissions, the Webhook responds with an AdmissionResponse with the allowed field set to false, causing the API Server to reject the request before it reaches the Operator.
Figure 11: Simplified Admission WebHook Example For Mitigating gateway-operator Vulnerability
We implemented a Validating Admission Webhook with Kubebuilder [kubebuilder]. Figure 11 illustrates a simplified version of our Webhook to mitigate the cross-ns secret reference vulnerability of the Kong gateway-operator. The Line 3-4 extracts the raw resource and raw request. The Line 6 gets the namespace of the referenced secret from the raw resource. The Line 7-16 crafts a SubjectAccessReview query, which checks whether the requesting user has permission to Get Secrets in the referenced namespace. The Webhook issues the query at Line 18, checks the result at Line 19, and rejects the unauthorized user at Line 20. With the scaffolding generated by Kubebuilder, tens of lines of code are required to implement a basic mitigation. And developers or cluster administrators can further implement other finer-grained validations in complex scenarios if needed.
TABLE VII: Validating Admission Webhook Mitigation Evaluation
Operator Mitigated Before (ms) After (ms)
Operator X 95.7 99.3 (+3.6)
gateway-operator 94.5 97.0 (+2.5)
baremetal-operator 91.6 94.5 (+2.9)
observability-operator 96.5 101.2 (+4.7)
tempo-operator 106.3 113.7 (+7.4)
ais-k8s 101.7 105.2 (+3.5)
Average 97.7 101.8 (+4.1)
We evaluate its effectiveness and overhead on the vulnerable Operators listed in Table VII. The evaluation is conducted on a Kubernetes cluster with 1 control plane node and 3 worker nodes. Each node is equipped with 8 cores of Intel E5-2680 v4 CPU (2.40GHz) and 32GB RAM. To evaluate the effectiveness, we created two ServiceAccounts in the cluster: one with cluster-level permissions and another limited to a single namespace. Before deploying the mitigation, both ServiceAccounts were able to invoke all listed vulnerable Operators to perform cross-ns references. After deploying the mitigation, only the ServiceAccount with legitimate cross-ns permissions could invoke the affected Operators, while the ServiceAccount with single-ns permissions was denied in all cases. This demonstrates that the mitigation successfully enforces namespace boundaries and prevents unauthorized privilege escalation.
To evaluate the overhead, since the Webhook only works at admission time and operates independently before the Operator logic, we measured the time between the point at which a new resource request was initiated and the point the Kubernetes API Server accepted the resource. For each test case, we repeated it 10 times and calculated the average overhead. Experimental results show an average increase of 4.1 ms after applying the mitigation. Given that the added overhead is a one-time overhead for a single resource request and Operator logic typically takes seconds to minutes, this added delay is negligible in practice.
To facilitate the mitigation, we open-sourced our Webhook for reference and a lightweight automatic Webhook generator, enabling quick adoption for practitioners.
Difference with Related Kubernetes Attacks. The cross-ns attack is fundamentally different from existing attacks [takeover, k8sblackhat, podescape] regarding threat models, root causes, and defenses.
For threat models, as illustrated in Section III, the precondition of launching existing attacks is to assume attackers have compromised containers (e.g., get a shell) before conducting overprivilege attacks, leaving critical gaps in how to compromise containers. In contrast, we presented a practical threat model without such strong assumptions, detailing two practical strategies to exploit Operators from scratch.
For root causes, existing attacks stem from granting applications unused permissions. Once attackers compromise applications (e.g., gain shells), they can conduct existing attacks by using over-granted privileges in follow-up actions. Thus, they [takeover, k8sblackhat, podescape, redhatreport, operatorsynk, operatorkubeops] suggest the Principle of Least Privilege (PoLP). Since only removing unused permissions, PoLP will not affect their functionality and is an ideal defense.
In contrast, our attack arises even when Operators follow PoLP. Supposing an Operator reading user input to get Secrets in a specified namespace (rather than assuming gain shells), it typically implies the Operator needs secret information for its operation, thus requiring certain privileges to function correctly. Removing such privileges would break the Operator’s functionality, even for legitimate intra-ns references. In other words, our vulnerability stems from insecure use of legitimate privilege, which existing defenses cannot address. An ideal defense should preserve necessary privileges while mitigating vulnerabilities; therefore, we propose new mitigations.
More Cross-Tenant Attacks. Similar cross-tenant attacks can also happen in various cloud scenarios beyond Kubernetes. Security researchers have discovered multiple services provided by Microsoft Azure [ChaosDB, extrareplica, synapse, ACSESSED, entraid], AWS [awsecr, appsync], Google Cloud Platform [gcpbigquery, gcplookstudio], and Oracle [attachme] were vulnerable to cross-tenant attacks, allowing attackers to manipulate unauthorized resources of other tenants. Our work further extends the landscape to Kubernetes and emphasizes cross-tenant risks in modern cloud architectures.
VII Related Work
@@
Kubernetes Operators and Controllers. Existing research on Kubernetes operators and controllers focuses on functional bugs instead of attacks. Acto [acto] proposes an automatic end-to-end testing technique for validating the operational correctness of Kubernetes Operators. Acto continuously generates desired state declarations and verifies whether the Operator correctly reconciles the system to those states. Sieve [sieve] presents an automatic reliability testing framework for cluster-management controllers. By injecting faults, Sieve uncovers deep semantic bugs by observing how controllers behave under fault conditions they are expected to tolerate. Anvil [anvil] presents the formal verification framework for Kubernetes controllers via TLA-style temporal reasoning, validating whether controllers eventually bring the cluster to the desired state and maintain it. Kivi [kivi] verifies Kubernetes controllers and their configurations by modeling controller behaviors and checking for violations of user-defined intent properties using model checking. It detects issues like imbalance and lifecycle bugs, focusing on functional correctness. operatorbug systematically summarize historical functional bugs of Operator. Red Hat [operatorsecpractice], Synk [operatorsynk], and KubeOps [operatorkubeops] suggest several good practices for Kubernetes Operators. However, their core guidance, limiting RBAC scope, is a general Kubernetes practice, not specific to Operator, and does not mitigate our proposed attack.
@@
To the best of our knowledge, our work is the first comprehensive study on Kubernetes Operator attacks.
@@
Kubernetes Security. In terms of attack and exploitation techniques, MITRE [mitrematrix] and Microsoft [msmatrix] summarize tactics to compromise containers and container orchestration systems like Kubernetes. kubedevopsprivesc investigate privilege escalation scenarios for DevOps pipelines on Kubernetes. ziyi present cross-container attacks on Kubernetes with eBPF. paced design Cross-Linux-Namespace defenses to detect container escape attacks. honeypot set up honeypots on Kubernetes and analyzes the attacks towards containers and container orchestration systems. coresidency present a co-residency attack towards container orchestration systems. qingyang comprehensively analyze 30 vulnerabilities in Kubernetes stacks. k8sblackhat, podkubecon identify the threat of trampoline pods which can be leveraged to gain escalated privileges. However, these works have not revealed or addressed the security issues brought about by Kubernetes operators.
@@
Kubernetes offers extensive configuration options for managing applications, including access controls and specifying security contexts. Any misconfigurations can lead to severe security vulnerabilities. Thus, another theme of Kubernetes security research is eliminating misconfiguration. Shamim et al. [xipractice, configurationsurvey, mitigatemisconfig] systematically reveal the risks of misconfiguration regarding best practice. kubemisconfigmeasure design static analysis tools and conduct a large-scale empirical study on Kubernetes manifests, revealing the landscape of misconfiguration. kgsecconfig leverage knowledge graphs to detect and mitigate Kubernetes misconfiguration. Recent work takeover identify the security risk of excessive Kubernetes RBAC permissions, which may lead to whole cluster takeover. EPScan [epscan] follows up on the research and designs systems to automatically minimize RBAC permissions. The industry also presents numerous tools for Kubernetes security, including Trivy [trivy], Kubescape [kubescape], KubeSec [kubesec], KubeArmor [kubearmor], Open Policy Agent [opa], and Kyverno [kyverno], providing functions like misconfiguration detection and runtime policy enforcement.
@@
While the existing works try to address misconfiguration and achieve the Principle of Least Privilege (PoLP) for applications, the vulnerability we present is not simply misconfigurations or violations of PoLP. They arise from inherent flaws in how Operators process user-controlled resources. These vulnerabilities exist even when the permissions of Operators are minimal, highlighting a deeper design-level security gap in the Operator model itself.
VIII Conclusion